OpenStudy (anonymous):

A marble is drawn at random from a bag that contains 3 green, 4 yellow, 5 blue, and 8 pink marbles. What is the probability that the marble drawn is green or blue?

OpenStudy (anonymous):

wouldn't that be green plus blue marbles / total marbles?

OpenStudy (anonymous):

3+5/3+4+5+8 = 0,4 = 40%

OpenStudy (anonymous):

2/5

OpenStudy (anonymous):

p(green) =3/20 , P(blue)=5/20 =1/4 , ===> P(green U blue) = (3/20)+(5/20)=8/20 = 2/5

OpenStudy (anonymous):

sample space=3+4+5+8=20 probability of drawing green= 3/20 probability of drawing blue= 5/20 the two events are independent and mutually exclusive so there is no intersections therefore probability of drawing blue or green =3/20 + 5/20 =8/20=2/5

OpenStudy (anonymous):

im sure with my answer...hehehheheh....100 percent probability